OpenStudy (falling_in_katt):

Simply where possible. \[(r ^{-3}s ^{2}t)^{4}\]

OpenStudy (zzr0ck3r):

Hint: \((ab)^c=a^cb^c\) \((a^b)^c=a^{bc}\)

OpenStudy (falling_in_katt):

\[rst ^{-4}?\]

OpenStudy (zzr0ck3r):

\((r ^{-3}s ^{2}t)^{4}= (r^{-3})^4(s^2)^4 t^4=r^{-3*4}s^{2*4}t^4\)